OpenStudy (kitten_is_back):

Starting at the bottom of the fire escape ladder on the outside of her building, Jiang Li climbed 63 steps to a friend’s apartment. She then descended 42 steps to retrieve her cat and then ascended 15 steps, coming to rest outside of her apartment’s window. On what step is Jiang Li standing? Step

OpenStudy (anonymous):

63 - 42 = 21 + 15 = 36 step